# DREAM: Dynamic Red-teaming across Environments for AI Models

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used in agentic systems, where their interactions with diverse tools and environments create complex, multi-stage safety challenges. However, existing benchmarks mostly rely on static, single-turn assessments that miss vulnerabilities from adaptive, long-chain attacks. To fill this gap, we introduce DREAM, a framework for systematic evaluation of LLM agents against dynamic, multi-stage attacks. At its core, DREAM uses a Cross-Environment Adversarial Knowledge Graph (CE-AKG) to maintain stateful, cross-domain understanding of vulnerabilities. This graph guides a Contextualized Guided Policy Search (C-GPS) algorithm that dynamically constructs attack chains from a knowledge base of 1,986 atomic actions across 349 distinct digital environments. Our evaluation of 12 leading LLM agents reveals a critical vulnerability: these attack chains succeed in over 70% of cases for most models, showing the power of stateful, cross-environment exploits. Through analysis of these failures, we identify two key weaknesses in current agents: contextual fragility, where safety behaviors fail to transfer across environments, and an inability to track long-term malicious intent. Our findings also show that traditional safety measures, such as initial defense prompts, are largely ineffective against attacks that build context over multiple interactions. To advance agent safety research, we release DREAM as a tool for evaluating vulnerabilities and developing more robust defenses.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) [chowdhery2023palm, brown2020language] have greatly improved AI capabilities, leading to their widespread adoption in critical areas like healthcare [clusmann2023future], finance [wu2023bloomberggpt, li2023large], and personal assistants [Gemini_Assistant]. The rise of conversational AI platforms like ChatGPT [OpenAI_ChatGPT], Gemini [Gemini], and Claude [Claude] has brought these capabilities to millions of users. Building on these foundation models [achiam2023gpt, chen2025janus, guo2025deepseek, qin2025ui], a new generation of autonomous AI agents is emerging [masterman2024landscape, cheng2024exploring], capable of navigating digital environments, managing tools and APIs, and executing complex multi-step workflows. While these capabilities present immense opportunities, they also introduce new security challenges that require urgent attention.
Figure 1: Comparison between Traditional Benchmarks and Our DREAM. Left: Traditional agent-safety benchmarks evaluate a model within a single environment using static, template-based attacks, offering limited coverage of realistic adversarial behavior. Right: Our proposed DREAM introduces a multi-agent attacker to evaluate models across diverse environments. A centralized Conductor performs cross-environment reasoning, while the Rater and Sandbox jointly evaluate and update attack states. This design enables dynamic, adaptive, multi-environment attack chains, uncovering vulnerabilities that static and single-environment tests fail to reveal.
Current safety evaluation methods [li2024salad, liu2024mmsafetybench, zeng2024air, chao2024jailbreakbench, yu2023gptfuzzer], however, are struggling to keep up with these advancements. Traditional red-teaming and safety benchmarks primarily focus on isolated tasks and assume a single-environment context. These evaluations, although useful, fail to capture how adversaries could exploit chains of seemingly benign interactions across different environments to achieve malicious objectives.
Real-world attacks expose the limitations of single-turn evaluations [ha2025one, li2025beyond, kumar2023certifying, perez2022red]. These attacks are dynamic and adaptive, often involving multi-step exploit chains where a small leak in one step leads to a critical breach in another. Long-chain attacks highlight two fundamental issues: agents’ contextual fragility, where safety mechanisms fail to transfer across environments, and their inability to track long-term malicious intent across multiple interactions.
The risk of multi-step attacks is not theoretical. Previous studies [li2023multi, paulus2024advprompter] have shown that agents can be manipulated into undesirable behaviors through carefully crafted prompt sequences. For example, seemingly harmless initial prompts can set up contexts that later enable harmful outputs. Moreover, safety protocols trained in one context can be bypassed in new environments, underscoring the challenge of contextual fragility. These studies reveal a critical gap: while individual attack vectors [yin2025safeagentbenchbenchmarksafetask, greshake2023not, jiang2024artprompt] have been documented, a systematic approach for evaluating agents’ robustness across extended, multi-context interactions is still lacking.
To address this gap, we introduce DREAM (Dynamic Red-teaming across Environments for AI Models), an automated framework for generating and evaluating dynamic, cross-environment attacks against AI agents, as shown in Figure 1. Unlike traditional methods that test static attack vectors, DREAM simulates a persistent adversary capable of conducting multi-step, adaptive attack campaigns. The system operates through an attack chain driven by an adversarial (red team) agent, which continuously refines its tactics based on accumulated information and responses to maximize the likelihood of successful exploitation. We investigate the following research questions: ❶ How can we automatically generate realistic, multi-step attack chains that simulate sophisticated adversarial behavior against agentic AI systems? ❷ Can cross-environment information pivoting reveal vulnerabilities that remain hidden in isolated, single-environment evaluations? ❸ What emergent vulnerabilities in agentic systems are missed by single-turn testing but exposed by multi-turn interactions? ❹ How do different advanced LLM agents compare in terms of their robustness against dynamic, adaptive attack chains?
@@
A key innovation of DREAM is its ability to automatically discover and exploit multi-stage, cross-environment vulnerabilities. This is achieved through an engine that mimics the strategic reasoning of a human adversary. At its core is a dynamic Cross‑Environment Adversarial Knowledge Graph (CE‑AKG), which fuses intelligence gathered from disparate environments into a unified, evolving world model. Guided by this global context, DREAM performs a Contextualized Guided Policy Search (C-GPS) to navigate the vast attack space. This closed-loop process allows the engine to opportunistically pivot across environments, construct long, causally-linked attack chains to amplify a domino effect, and adaptively re-plan by backtracking from failed attempts. This dynamic, stateful approach uncovers systemic risks that are invisible to static, single-turn evaluations.
Our Contributions. In developing and evaluating DREAM, our primary contributions include:
  • We present DREAM, an end-to-end framework for automated red teaming of LLM agents, capable of generating dynamic, multi-stage attack chains that span multiple digital environments.
  • We introduce the Cross‑Environment Adversarial Knowledge Graph (CE‑AKG), a novel, LLM-powered mechanism for stateful, cross-domain situational awareness. The CE-AKG enables an attacker agent to bridge information silos and leverage intelligence from one context to inform attacks in another.
  • We design the Contextualized Guided Policy Search (C-GPS) algorithm to operationalize the CE-AKG. This algorithm employs a strategic heuristic value function to construct long, causally-dependent attack sequences, effectively simulating the domino effect seen in sophisticated, real-world breaches.
  • We conduct a large-scale systematic evaluation of 12 state-of-the-art LLM agents, demonstrating DREAM’s effectiveness in uncovering emergent, long-chain vulnerabilities missed by conventional evaluations. Our findings reveal critical fragility patterns in modern agents, highlighting the urgent need for more robust, context-aware defense strategies.
Paper Organization. This research provides a more realistic and powerful tool for adversarial testing, helping to build the next generation of robust and secure AI agents. The remainder of this paper is organized as follows: § 3 reviews related work. § 4 details our proposed framework. § 5 presents the experimental setup and results, followed by a discussion in § 6 and our conclusion in § 7.
Figure 2: Overview of our DREAM framework. Left: Multi-Agent Atom Attack Generation, where Scout, Seeder, and Exploiter roles create diverse atom attacks for the Atom Attack Library. Center: The Cross-Environment Chain Attack powered by Contextualized Guided Policy Search (C-GPS), with the Conductor agent dynamically planning attack paths through retrieval, clustering, action selection, and failure-aware backtracking. Right: A case study illustrating a three-step attack, leveraging information from one environment to initiate a cross-environment exploit, facilitated by the Cross-Environment Adversarial Knowledge Graph (CE-AKG) to trigger a domino effect and uncover systemic vulnerabilities.
2 Preliminary Study
To evaluate current safety approaches, we ran a pilot study using gemini-2.5-flash as a representative model. Our analysis tested agent robustness across three key areas: the effects of longer attack sequences, the dangers of switching between environments, and how well static defenses hold up.
Our analysis reveals a nonlinear relationship between attack chain length and success. While individual attacks often fail to break through safety measures, longer sequences produce dramatically better results. We call this the “domino effect,” showing that single-turn evaluations seriously underestimate threats from attackers who maintain malicious intent over time. Breaking down harmful goals into apparently harmless steps lets attackers bypass defenses that work fine in isolation.
Beyond chain length, the assumption that environments are isolated also causes problems. Our results show that switching between different digital domains greatly increases attack success. Agents that seem robust in one environment often show “contextual fragility” when information crosses between contexts. This suggests vulnerabilities often emerge from cross-domain interactions, not just isolated flaws.
Most importantly, DREAM moves away from static attack templates toward dynamic, feedback-based testing. Unlike fixed benchmarks with predetermined attacks, our framework lets attackers adapt strategies based on model responses. This dynamic approach lets researchers adjust complexity through parameters like chain length, testing both minimal safety under simple attacks and maximum vulnerability under sustained pressure. As a result, our variable approach resists data contamination and overfitting, solving the common problem of benchmark leakage.
3 Related Work
In this section, we review the literature that forms the foundation for our work. We first discuss recent progress in traditional LLM safety evaluation, which focuses on the models themselves. Then, we examine the emerging field of agent safety evaluation, which addresses the unique challenges of LLM-powered autonomous systems.
3.1 LLM Safety Evaluation
The widespread adoption of large language models (LLMs) has sparked significant research into their safety risks [bengio2024managing]. One line of work develops static safety benchmarks to evaluate model behavior [sun2023safety, xu2023cvalues, cui2023fft, zhang2024safetybench]. These benchmarks use comprehensive classification systems and various evaluation methods. Common approaches include direct prompting and multiple-choice questions. However, even well-aligned LLMs remain vulnerable to jailbreak attacks. As a result, another line of research incorporates adversarial techniques into evaluations [zou2023universal, li2024salad, mazeika2024harmbench, huang2024flames, chao2024jailbreakbench, xie2024sorry, dong2024attacks]. These techniques test whether models can be tricked into generating unsafe content.
3.2 Agent Safety Evaluation
Agents have shown impressive abilities to solve complex tasks through reasoning and planning. They interact effectively with real-world environments [qin2023toolllm]. However, their safety concerns go beyond those of traditional LLMs. LLMs mainly risk generating unsafe content. In contrast, agents can perform potentially harmful actions in interactive environments. Yin et al. [yin2024safeagentbench] explored risks with embodied LLM agents in real-world scenarios. Lee et al. [lee2024mobilesafetybench] introduced MobileSafetyBench for evaluating device-control agents. Their benchmark uses Android emulators to simulate realistic mobile environments. Recently, research has focused on agents that interact with external tools. These agents work across diverse domains like banking, travel, and online shopping [yuan2024r, debenedetti2024agentdojo, xiang2024guardagent, ruan2023identifying, ye2024toolsword, zhan2024injecagent]. Despite these advances, current benchmarks remain limited in scope. We need a more comprehensive benchmark. It should cover a wider variety of environments. It should also provide systematic coverage of diverse risk categories. Finally, it must address various failure modes in tool-augmented agents.
4 Methodology
4.1 Formalizing Attack Chain Generation
We model the generation of an attack chain $A=(a_{1},\dots,a_{T})$ as a policy search problem within a Partially Observable Markov Decision Process (PO-MDP), defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{O},P,R,\gamma\rangle$.
Let $\mathcal{S}$ be the latent state space of the target system, and let $b_{t}\in\mathcal{B}(\mathcal{S})$ be the belief state of the Conductor agent at timestep $t$. The action space $\mathcal{A}$ consists of a discrete set of atomic attacks. After executing action $a_{t}\in\mathcal{A}$, the Conductor receives a new observation $o_{t+1}\in\mathcal{O}$ (a natural language response) and updates its belief according to the deterministic function $\tau$:
$b_{t+1}=\tau(b_{t},a_{t},o_{t+1}),$ (1)
with the objective of finding an optimal policy $\pi:\mathcal{B}(\mathcal{S})\to\mathcal{A}$ that maximizes the expected cumulative reward:
$E_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}R(b_{t},a_{t})\right].$ (2)
Due to the high dimensionality and intractability of this problem, we approximate the optimal policy $\pi^{*}$ with a heuristic policy, $\pi_{\text{cond}}$, which selects actions by maximizing a value function $V(b_{t},a)$ over a dynamically generated candidate set $\mathcal{C}_{t}\subseteq\mathcal{A}$:

### v2

# DREAM: Dynamic Red-teaming for Evaluating Agentic Multi-Environment Security

@@ Abstract
As Large Language Models (LLMs) evolve from passive text generators into autonomous agents, their safety profile becomes inextricably linked to their continuous interaction with external tools and dynamic environments. Current safety benchmarks, however, remain largely confined to static, single-turn assessments, failing to capture the risks emerging from the temporal and adaptive nature of agentic behaviors. In these interactive settings, adversaries can leverage environment feedback to dynamically calibrate their strategies, or hide malicious intent within long-chain sequences of seemingly innocuous tool calls. To address these sophisticated threats, we present DREAM, a systematic framework for evaluating the interaction-level safety of LLM agents through dynamic, multi-turn adversarial simulations. At the core of DREAM lies the Cross-Environment Adversarial Knowledge Graph (CE-AKG), a novel abstraction that formalizes the orchestration of multi-stage exploits. By treating existing single-turn static attacks as fundamental “atomic actions,” CE-AKG strategically leverages a Contextualized Guided Policy Search (C-GPS) to assemble atomic actions into comprehensive, long-chain attack trajectories. A comprehensive evaluation of 12 sota LLM agents reveals a stark reality: over 68% of long-chain exploits successfully bypass existing defenses, exposing a fundamental gap in stateful, cross-environment security.
@@ 1 Introduction
The evolution of LLMs [5, 2, 42, 52, 6] into autonomous agents has shifted the security paradigm from static prompt inspection to managing stateful interaction trajectories. Unlike passive generators, these systems’ continuous tool use expands the attack surface, enabling adaptive “long-chain” exploits [22, 4, 35, 34, 39] where individually benign actions cumulatively bypass defenses. The urgency of this threat is highlighted by emerging architectures like OpenClaw (formerly Clawdbot) [31]. By bridging external messaging platforms with local operating systems, such systems exemplify the precise cross-environment “contextual fragility” DREAM is designed to simulate—a vulnerability chain where a remote prompt triggers local execution, effectively evading traditional static benchmarks.
Figure 1: DREAM vs. Traditional Benchmarks. Left: Static, single-environment evaluation with fixed attack templates. Right: DREAM’s multi-agent system (Conductor, Rater, Sandbox) enables dynamic cross-environment reasoning and discovery of long-chain exploits.
Despite the high stakes, current safety evaluation methods [18, 20, 48, 3, 44, 21, 38] are struggling to keep up with these advancements. Traditional red-teaming and safety benchmarks remain largely confined to stateless and static assessments [11, 19, 15, 26]. Such evaluations, while useful for catching explicit harms, fail to account for the dynamic and adaptive nature of real-world exploits. In practice, sophisticated adversaries do not rely on a single malicious prompt, but instead orchestrate chains of seemingly benign interactions across disparate environments to bypass filters and achieve a critical breach. This reveals a fundamental gap: existing benchmarks [47, 29, 51, 46] predominantly operate on a stateless paradigm, treating safety as an isolated invariant rather than a stateful interaction trajectory. Consequently, they fail to correlate fragmented signals into a coherent malicious intent across extended, cross-environment sequences.
To address this systemic vulnerability, we introduce DREAM (Dynamic Red-teaming for Evaluating Agentic Multi-Enviroment Security), an automated framework for generating and evaluating multi-step, cross-environment attacks (Figure 1). Unlike conventional methods [17, 25, 10, 14] that test static attack vectors, DREAM models a persistent adversary capable of conducting adaptive campaigns. By operating as a closed-loop system, the framework iteratively calibrates its tactics, fusing historical observations with immediate environmental responses. This approach allows us to probe the “contextual fragility” of agents, a phenomenon in which safety mechanisms effective in one environment fail to generalize when information is pivoted across another.
@@
The core innovation of DREAM operationalizes human-like strategic reasoning through a Cross-Environment Adversarial Knowledge Graph (CE-AKG). The CE-AKG serves as a formal abstraction that fuses intelligence from isolated environments into a unified, evolving world model. Guided by this global context, DREAM leverages a Contextualized Guided Policy Search (C-GPS) to navigate the vast attack space. This mechanism enables the engine to exploit cross-environment “pivot points,” orchestrate causal chains to trigger a domino effect, and employ failure-aware backtracking.
To empirically demonstrate the obsolescence of existing static benchmarks, we utilize DREAM to conduct a large-scale systematic evaluation of 12 state-of-the-art LLM agents. Rather than merely probing for potential risks, our evaluation is designed to rigorously stress-test the prevailing “stateless” safety paradigm against dynamic, cross-environment attack chains. By subjecting these agents to adaptive adversarial campaigns that mimic sophisticated human strategies, we aim to quantify the “blind spots” inherent in current evaluations and expose the systemic inability of modern architectures to withstand stateful exploitation.
Our evaluation reveals an alarming reality: over 68% of long-chain exploits successfully bypass existing defenses. More critically, attacks utilizing 5-step chains across 5 environments expose vulnerability severities that are 4.8 times higher than those estimated by static baselines. This disparity highlights that current safeguards are largely “context-unaware,” focusing on individual inputs while remaining blind to the stateful orchestration of an attack. In summary, we present the following contributions:
  • We formalize a shift in agent evaluation from stateless, atomic testing to stateful trajectory auditing. By defining the interplay between temporal sequences and cross-environment dependencies, we establish a new paradigm for assessing the logical resilience of LLM agents under sustained adversarial pressure.
  • We introduce DREAM, a closed-loop red-teaming framework powered by the Cross-Environment Adversarial Knowledge Graph (CE-AKG). This technical innovation enables the autonomous synthesis of complex attack chains by fusing fragmented environment feedback into a unified strategic world model, allowing for adaptive re-planning that mimics human adversarial reasoning.
  • A large-scale evaluation of 12 SOTA agents reveals that current defenses are largely ineffective against long-chain exploits, uncovering critical fragility patterns that highlight the urgent need for context-aware safeguards.
Roadmap.
The paper is organized as follows: Section 2 formalizes the theoretical problem of dynamic adversarial interactions. Section 3 details the proposed DREAM framework. Section 4 presents the experimental results, and analysis. Section 5 concludes the paper. To provide a comprehensive view and due to page limitation, we include the related work, Discussions on defenses and real-world implications, statistical significance tests, qualitative case studies, and formal metric definitions in the Appendix.
2 Methodologies
This section formalizes the operational paradigm of tool-augmented agents, contrasting static, single-turn interactions with dynamic, multi-environment workflows. This formalization highlights the limitations of stateless evaluation and establishes the basis for DREAM. Table 1 summarizes the key notations used throughout.
Table 1: Mathematical notation and key variables in DREAM.
Symbol Description
PO-MDP Formulation (Sec. 2.3)
     $\mathcal{S},\mathcal{A},\mathcal{O}$ Latent state, Action, and Observation spaces
     $P,R,\gamma$ Transition function, Reward function, and Discount factor
     $b_{t}$ Belief state at timestep $t$ (Implemented via CE-AKG)
     $\tau$ Belief update function (Information Fusion)
     $\pi^{*}$ Optimal policy approximating the vulnerability Upper Bound
Theoretical Analysis (Sec. 2.4)
     $\mathcal{T}$ Dynamic attack trajectory, $\mathcal{T}=(b_{0},a_{0},o_{1},\dots)$
     $\rho$ Attack Intent Density ($\mathcal{I}(G)/\Delta T$)
     $\mathcal{H}(b_{t})$ Entropy of belief state (Epistemic Uncertainty)
     $K_{accum}$ Accumulated Knowledge Context (Entropy Reduction)
     $\Omega$ Defense Entropy Barrier (Phase Transition Threshold)
Implementation & Metrics (Sec. 3)
     $a_{t},o_{t}$ Action executed and Observation received at step $t$
     $\mathcal{C}_{t}$ Candidate action set generated by C-GPS
     $R(b_{t},a_{t})$ Immediate reward for action $a_{t}$
     $\text{Score}(\mathcal{T})$ Final discounted cumulative reward of trajectory $\mathcal{T}$
2.1 The Target: Tool-Augmented Agent Systems
Modern autonomous agents have evolved from passive text generators into sophisticated tool-augmented systems capable of interacting with multiple heterogeneous environments. Through standardized interfaces [27], a single Large Language Model (LLM) acts as a central controller, coordinating actions across a cluster of diverse domains $\mathcal{E}=\{E_{1},E_{2},\dots,E_{k}\}$ within a unified context. In this paradigm, the agent receives user instructions and autonomously invokes domain-specific tools (e.g., database queries, API calls, file manipulations) to execute complex tasks. Consequently, the safety surface of such an agent is not defined by a single input-output pair, but by the security of the entire interaction history across $\mathcal{E}$.
2.2 The Threat Model: Dynamic Attack Trajectories
Existing benchmarks predominantly evaluate safety using static, single-turn prompts, effectively modeling an attack as an isolated pair $(a,o)$. However, such static snapshots fail to capture the complexity of real-world exploitation in agentic systems. To bridge this gap, we redefine the attack not as a fixed input, but as a dynamic trajectory $\mathcal{T}$ over a time horizon $T$:
$\mathcal{T}=(b_{0},a_{0},o_{1},b_{1},\dots,a_{T-1},o_{T}),$ (1)
where $b_{t}$ represents the attacker’s accumulated knowledge state and $a_{t}$ is an action dependent on that knowledge. In this context, the ultimate goal of the threat model is not merely to test for the presence of known vulnerabilities, but to discover the optimal trajectory $\mathcal{T}^{*}$ that maximizes the exposed security risk. Consequently, this shifts the evaluation paradigm from compliance checking, which sets a risk lower bound, to worst-case analysis, which exposes the vulnerability upper bound by discovering the maximum damage an adaptive adversary can inflict.
2.3 Problem Formulation: Adversarial Interaction as PO-MDP
As highlighted in Section A, achieving the adversarial objective defined above requires capabilities beyond current stateless evaluations. To effectively discover the optimal trajectory $\mathcal{T}^{*}$, an evaluation framework must address three critical requirements:
  • Cross-Environment Pivoting: The ability to bridge information silos, leveraging data found in environment $E_{i}$ to compromise $E_{j}$.
  • The Domino Effect: The capacity to model causal dependencies, where the feasibility of a high-impact attack strictly depends on the accumulated success of prior benign actions.
  • Dynamic Adaptation: The capability to transition from estimating the vulnerability lower bound to approximating the vulnerability upper bound via adaptive, multi-turn policy search.
To systematically operationalize these requirements, we model the adversarial interaction as a Partially Observable Markov Decision Process (PO-MDP), defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{O},P,R,\gamma\rangle$. In this framework, the attacker selects an action from the space $\mathcal{A}$, which triggers the environment to transition according to the hidden dynamics $P$ and emit an observation from the space $\mathcal{O}$.
Latent State & Belief State. Let $\mathcal{S}$ denote the true, latent state of the target system including hidden files, internal agent logic, and security configurations. Since the attacker cannot observe $\mathcal{S}$ directly, they maintain a Belief State $b_{t}$, which represents the attacker’s current knowledge graph of the system and addresses the challenge of partial observability.
Belief Dynamics. The core mechanisms of the “Domino Effect” and “Cross-Environment Pivoting” are mathematically captured by the belief update function $\tau$. At each step, the attacker refines their knowledge based on the observation $o_{t+1}$ triggered by action $a_{t}$:
$b_{t+1}=\tau(b_{t},a_{t},o_{t+1}).$ (2)
This recursive update function $\tau$ ensures that $b_{t+1}$ aggregates context (e.g., credentials from $E_{i}$) unlocked by previous actions, making it available for subsequent attacks in $E_{j}$. In our framework, this is concretely implemented via the Cross-Environment Adversarial Knowledge Graph (CE-AKG).
Optimization Goal. This optimization formulation provides a unified theoretical basis for assessing security bounds through trajectory length and policy adaptivity. We seek an optimal policy $\pi^{*}$ that maximizes the expected cumulative reward, guided by the reward function $R$ and the discount factor $\gamma$:
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\mathcal{T}\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}R(b_{t},a_{t})\right].$ (3)
Here, $R(b_{t},a_{t})$ quantifies the immediate utility of an attack step, while $\gamma$ balances immediate versus long-term gains. By solving for $\pi^{*}$, this formulation allows us to measure the full spectrum of risk. A restrictive policy (e.g., starting from $T=1$) establishes the vulnerability lower bound, while expanding the search horizon $T$ approximates the vulnerability upper bound via our C-GPS.
2.4 Theoretical Justification: Entropy and Phase Transitions
While the PO-MDP framework provides the structural basis for trajectory optimization, we now analytically demonstrate why dynamic trajectories ($\Delta T>1$) can strictly outperform static baselines ($\Delta T=1$) in bypassing defense thresholds. We model this as an information-theoretic state transition of the belief state $b_{t}$. The Detectability Bound of Static Attacks. Let $\mathcal{I}(G)$ denote the minimum information payload (in bits) required to trigger a specific malicious goal $G$. We define the Attack Intent Density $\rho$ as:
$\rho=\frac{\mathcal{I}(G)}{\Delta T}.$ (4)
In static benchmarks where $\Delta T=1$, the intent density is maximized ($\rho_{max}=\mathcal{I}(G)$). This high-density payload forces the exposure of high-entropy malicious features that likely exceed the safety detection threshold $\theta_{safe}$ of stateless filters (i.e., $\rho>\theta_{safe}$), leading to immediate interception. This explains why static testing often yields a falsely optimistic lower bound of vulnerability.
DREAM: Epistemic Uncertainty Reduction. In contrast, DREAM distributes $\mathcal{I}(G)$ across a trajectory $\mathcal{T}$. In the early stages ($t<T$), the Conductor prioritizes Information Gain to reduce epistemic uncertainty. Let $\mathcal{H}(b_{t})$ be the entropy of the attacker’s belief state at step $t$. The update process $b_{t}=\tau(b_{t-1},a_{t-1},o_{t})$ results in an entropy reduction:
$\mathcal{H}(b_{t})=\mathcal{H}(b_{t-1})-I(a_{t};\mathcal{O}_{t}),$ (5)
where $I(a_{t};\mathcal{O}_{t})$ is the mutual information extracted from observation $\mathcal{O}_{t}$. Consequently, the Accumulated Knowledge $K_{accum}$ contained within $b_{t}$ grows progressively:
$K_{accum}(t)=\mathcal{H}(b_{0})-\mathcal{H}(b_{t})=\sum_{i=1}^{t}I(a_{i};\mathcal{O}_{i}).$ (6)
This allows the attacker to maintain a low local intent density ($\rho_{t}\ll\theta_{safe}$) while $b_{t}$ silently accumulates the necessary context.
The Phase Transition of “Thunderous Strike”. To model the non-linear jump from stealthy preparation to successful exploitation, we introduce a Knowledge-driven Phase Transition function using the Hill Equation:
$P_{success}(b_{t})=\frac{(K_{accum}(t)\cdot\omega)^{k}}{\Omega^{k}+(K_{accum}(t)\cdot\omega)^{k}}.$ (7)
Here, $\Omega$ represents the Defense Entropy Barrier (the information complexity of the safety mechanism), $\omega$ is the strategic efficiency coefficient, and $k$ is the cooperativity coefficient. For $t<T$, where $K_{accum}\cdot\omega<\Omega$, the probability $P_{success}\approx 0$, creating a “Safety Illusion.” However, once the knowledge in $b_{t}$ exceeds the barrier ($K_{accum}\cdot\omega>\Omega$), $P_{success}$ undergoes a catastrophic leap. This mathematically characterizes the Contextual Fragility of agents: they remain robust until the precise moment the adversarial context in $b_{t}$ is fully assembled.
3 The DREAM Framework
Figure 2: The framework of DREAM. Left: Specialized agents generate diverse atomic attacks for the library. Center: The Conductor orchestrates cross-environment trajectories using C-GPS, featuring adaptive action selection and backtracking. Right: A three-step attack sequence illustrating how the CE-AKG enables information pivoting to trigger a domino effect and reveal long-chain vulnerabilities.
3.1 System Instantiation: From Theory to Framework
To solve the PO-MDP optimization problem defined in Eq. (3) and operationalize the phase transition dynamics detailed in Section 2.4, DREAM instantiates the abstract mathematical components into a concrete, closed-loop adversarial framework. As illustrated in Figure 2, DREAM functions as a stateful search engine driven by an adversarial Conductor agent.
This architecture explicitly maps our theoretical formulation to system modules:
  • Belief State ($b_{t}$) $\to$ Unified Sandbox (CE-AKG): Implements the belief update function $\tau$ by fusing heterogeneous observations into a structured knowledge graph, effectively tracking the accumulated knowledge context $K_{accum}$.
  • Action Space ($\mathcal{A}$) $\to$ Atom Attack Library: Provides the discrete, low-intent-density primitives for trajectory construction.
  • Policy ($\pi^{*}$) $\to$ C-GPS Algorithm: Approximates the optimal adversarial policy through heuristic tree search and backtracking, guiding the Conductor to surmount the defense entropy barrier $\Omega$ and discover the vulnerability upper bound.
The following sections detail the construction of these components and the execution pipeline.
3.2 Building the Foundation: Action and State Spaces
Before presenting our planning algorithm, we detail the concrete instantiation of the action space $\mathcal{A}$ and state space representation that enable cross-environment attack reasoning.
3.2.1 Action Space: Multi-Agent Atom Attack Library
The action space $\mathcal{A}$ consists of atom attacks, structured primitives encoding all information needed for execution. Each atom $a\in\mathcal{A}$ contains: (1) a semantic description for retrieval, (2) a target environment $E\in\mathcal{E}$, (3) a parameterized prompt template, and (4) entity requirements specifying prerequisite information (e.g., user IDs, credentials) needed from the belief state.
Generation via Multi-Agent Roles. We systematically construct $\mathcal{A}$ by abstracting adversarial behaviors into three strategic roles inspired by red team operations:
  • Scout: Reconnaissance attacks that reduce uncertainty in $b_{t}$ by discovering new entities and system information (e.g., “enumerate user accounts”, “list accessible databases”).
  • Seeder: State manipulation attacks that alter $\mathcal{S}$ to create vulnerabilities for subsequent exploitation (e.g., “inject malicious configuration”, “plant backdoor credentials”).
  • Exploiter: Direct exploitation attacks that leverage current $b_{t}$ to achieve malicious objectives and earn high immediate rewards (e.g., “exfiltrate customer data”, “execute unauthorized transactions”).
LLMs are utilized with this multi-agent framework to generate attacks covering all three roles across each environment $E_{i}\in\mathcal{E}$. This semi-automated process yields a comprehensive library of 1,986 atom attacks spanning 349 distinct environments, forming a rich and challenging action space.
3.2.2 State Space: Cross-Environment Adversarial Knowledge Graph
A critical challenge in multi-environment attacks is overcoming information silos, where knowledge gained in one domain must inform actions in others. We address this through the Cross-Environment Adversarial Knowledge Graph (CE-AKG), which serves as the concrete implementation of the Conductor’s belief state $b_{t}$.
The Unified Sandbox Component. We realize the CE-AKG through the Unified Sandbox, an LLM-powered component that maintains $b_{t}$ as a dynamic JSON structure representing typed entities (user IDs, credentials, file paths, API keys, etc.). The Unified Sandbox implements the belief update function $\tau$ (Eq. (2)) via two key operations:
  • Information Fusion: After receiving observation $o_{t+1}$, the LLM parses the text, extracts newly discovered entities $\mathcal{K}_{t+1}$, and merges them into the existing belief state: $b_{t+1}=b_{t}\cup\mathcal{K}_{t+1}$. This ensures knowledge from any environment enriches the global state representation.
  • Context Provisioning: To instantiate the selected action $a_{t+1}$, the Unified Sandbox queries the CE-AKG based on the action’s specific entity requirements. It retrieves the relevant context $\text{ctx}_{t+1}$ from $b_{t+1}$ and instantiates the action’s prompt template, thereby equipping the attack with all previously acquired information regardless of source environment.
Enabling Cross-Domain Exploitation. This mechanism establishes an “information bridge” across security boundaries. For example: a customer_id discovered by a Scout in the MissionControl environment is fused into the CE-AKG and can later provision a targeted Exploiter attack in the RetailFraud environment. This cross-environment knowledge flow is essential for constructing logically coherent, multi-hop attack chains that span isolated domains.
3.2.3 Reward Function: Quantifying Exploitation Progress
To guide the Conductor’s policy, we define a reward function $R(b_{t},a_{t})$ that quantifies immediate exploitation utility. The core component is the Atomic Score, $\text{Score}(a)$, representing each attack’s intrinsic potential (success rate and impact), pre-computed via large-scale automated benchmarking.
Let $s(a_{t})\in\{0,1\}$ indicate whether action $a_{t}$ succeeded, which is determined by LLM-based evaluation of $o_{t+1}$). The immediate reward is:
$R(b_{t},a_{t})=s(a_{t})\cdot\text{Score}(a_{t})-(1-s(a_{t}))\cdot C_{\text{penalty}},$ (8)
where $C_{\text{penalty}}$ is a constant penalty for failures. This structure encourages successful, high-impact actions while penalizing dead-end paths, providing the foundation for strategic planning.
3.3 Contextualized Guided Policy Search (C-GPS)
To approximate the optimal policy $\pi^{*}$ and reveal the upper bound of system vulnerability, we propose the Contextualized Guided Policy Search (C-GPS). Unlike static templates that strictly adhere to pre-defined execution graphs, C-GPS orchestrates a dynamic tree search over the belief state space. It utilizes a heuristic value function to prune low-probability branches while employing backtracking to mitigate local optima. This ensures that if a viable high-severity attack chain exists within the horizon $T$, the Conductor maximizes the probability of discovering it.
3.3.1 The Planning Challenge
Directly optimizing the objective (Equation 3) faces two fundamental obstacles: (1) exponential search space, where the combination of $|\mathcal{A}|=1{,}986$ atoms and trajectory length $T\approx 10{-}20$, exhaustive search is infeasible, and (2) long-term credit assignment, as early reconnaissance actions may have zero immediate reward but unlock high-value exploits 10+ steps later. C-GPS addresses these challenges through heuristic value estimation combined with strategic candidate pruning.
3.3.2 Heuristic Value Function
We approximate the optimal policy via: